# Big bang nucleosynthesis with rapidly varying G

### v1

@@ II Models with Rapidly-Varying G
Step function models for $G$ have been explored in connection with inflation Ash and, more recently, as a possible resolution of the Hubble tension Marra ; Benevento ; Kable ; Alestas1 ; Alestas2 . Other observational consequences of a step function change in $G$ are explored in Refs. Alestas1 ; Alestas2 ; Paraskevas . Models of this kind are easily generated in the context of scalar-tensor theories with a step function $F(\phi)$ (see, e.g., Refs. Ash ; Paraskevas ), and they naturally asymptote to $G=G_{N}$, with $\dot{G}=0$, at late times.
@@ III Effects on Big Bang Nucleosynthesis
The weak reactions drop out of thermal equilibrium at $T\sim 1$ MeV, while free neutron decay continues until $T\sim 0.1$ MeV, when rapid fusion into heavier elements occurs. At this point, nearly all of the remaining neutrons become bound into ${}^{4}$e, with a small fraction incorporated into deuterium, tritium, and ${}^{3}$e. There is also production of ${}^{7}$i and ${}^{7}$e, with the latter decaying into the former via electron capture at the beginning of the recombination era Sunyaev . The element abundances produced in BBN depend on the baryon/photon ratio $\eta$, which can be independently determined from the CMB. We adopt a value of $\eta=6.1\times 10^{-10}$, consistent with results from Planck Ade . This value of $\eta$ yields predicted abundances of D and ${}^{4}$e consistent with observations. On the other hand, the ${}^{7}$i abundance predicted by BBN is well known to differ from observationally-inferred values, giving rise to the so-called “lithium problem” lithium . In what follows, we will consider only the ${}^{4}$e and deuterium production. Following standard practice, we give the primordial ${}^{4}$e abundance as a mass fraction, $Y_{p}$, and the deuterium abundance as a number density relative to hydrogen, (D/H).
When the change in $G$ is taken to be a constant during BBN, the effect on the ${}^{4}$e and deuterium abundances is well fit by FOYY
$Y_{p}=Y_{p0}(G/G_{N})^{0.357}$ (13)
@@
${\rm(D/H)}={\rm(D/H)}_{0}(G/G_{N})^{0.952}$ (14)
where the 00 subscript denotes the standard values of these abundances with an unmodified gravitational constant. Although both the ${}^{4}$e and deuterium abundances are increasing functions of $G$, this dependence arises for different reasons in these two cases. The ${}^{4}$e abundance depends mostly on the abundance of free neutrons when rapid fusion commences. An increase in $G$ causes the weak rates to freeze out at a higher temperature, when the equilibrium abundance of free neutrons is larger. Since nearly all of these neutrons end up in ${}^{4}$e, the result is a larger ${}^{4}$e abundance. However, changing $G$ also has a second effect on the ${}^{4}$e abundance. A larger $G$ corresponds to a smaller Hubble time at a given temperature. Thus, the temperature at which rapid fusion begins corresponds to a smaller time over which free neutron decay can occur, giving more neutrons to bind into ${}^{4}$e. This dual dependence on $G$ can be seen in Ref. Bambi , which presents response functions for the various element abundances as a function of a change in $G$ that is effectively a delta function of time. For $Y_{p}$, the response function is always positive, with dual peaks at $T\sim$ 1 MeV and $T\sim 0.1$ MeV, corresponding to the two effects noted above. Deuterium, on the other hand, is sensitive almost entirely to the value of $G$ when fusion begins. Given an arbitrarily long time, essentially all of the deuterium would fuse into heavier nuclei; the existence of primordial deuterium is a result of the relatively short time over which fusion can occur. Increasing $G$ decreases this time, resulting in less fusion of deuterium into heavier elements and a corresponding increase in (D/H). This is evident in the response function for deuterium given in Ref. Bambi . This response function is positive and sharply peaked near $T\sim 0.1$ MeV.
Figure 1: Change in the ${}^{4}$e mass fraction, $\Delta Y_{p}$, for a step-function variation in $G$ (Eqs. 7 - 9) as a function of $t_{i}$ and $t_{f}$, with $A=-0.1$ using (left) only the standard Friedman equation (1) and (right) the full Friedman equation (6) including the appropriate $\dot{G}/G$ terms.
@@
To calculate the primordial element abundances, we have modified PRyMordial, a Python code for BBN calculations primordial , incorporating Eq. (6) for the Friedman equation. To highlight the specific effects of rapid changes in $G$, we have also calculated the primordial element abundances using the standard Friedman equation (1) with time-varying $G$. We have chosen to present our results in terms of the change in the primordial ${}^{4}$e mass fraction, $\Delta Y_{p}$, and the change in the number density of deuterium relative to hydrogen, $\Delta$(D/H), due to the time variation in $G$, rather than giving the altered values of $Y_{p}$ and (D/H) themselves. This approach has the advantage that $\Delta Y_{p}$ and $\Delta$(D/H) are less sensitive to the input parameters such as the baryon-to-photon ratio, the neutron lifetime, and changes to the nuclear cross sections, because altering these parameters will tend to change both the standard-model and nonstandard values of these abundances in roughly the same way. Although we provide constraints on $\dot{G}/G$ for the rapidly-oscillating models, such constraints are subject to future changes in the observed element abundances, so we consider our results for $\Delta Y_{p}$ and $\Delta$(D/H) to be more interesting and fundamental.
@@
Consider first the step-function evolution for $G$ (Eqs. 7 - 9). Our results for the ${}^{4}$e abundances in this case with $A=-0.1$ and $A=+0.1$ are displayed in Figs. $1-2$. First consider the effect of simply altering $G$ in the standard Friedman equation without including the $\dot{G}/G$ terms. When $t_{i}$ is large, this is simply equivalent to BBN with a different (constant) value of $G$: here $G=(0.9)(G_{N})$ (Fig. 1, left) or $G=(1.1)(G_{N})$ (Fig. 2, left). As we have already noted, ${}^{4}$e is most sensitive to the value of $G$ at early times; for large $t_{i}$, we simply obtain a constant increase or decrease in the ${}^{4}$e abundance, given by the corresponding constant change in $G$ as in Eq. (13). However, when we use the full expression for the scale factor given by Eq. (6) (Figs. 1-2, right), we see that the $\dot{G}/G$ terms strongly modify the change in the primordial ${}^{4}$e production, producing more complex behavior. In particular, for $A<0$, we have $\dot{G}/G>0$, so the $\dot{G}/G$ terms increase the value of $\dot{a}/a$. This partially cancels the effect of the smaller initial value of $G$, producing a smaller decrease in $Y_{p}$. The opposite effect occurs when $A=+0.1$. In this case, $\dot{G}/G$ is negative, resulting in a decrease in $\dot{a}/a$ that partially cancels the effect of the initially larger value of $G$. Thus, in both cases, the effect of including the additional $\dot{G}/G$ terms in the expression for $\dot{a}/a$ in Eq. (6) is to reduce the overall effect of the change in $G$.
@@
Figure 3: Change in D/H for a step-function variation in $G$ (Eqs. 7 - 9) as a function of $t_{i}$ and $t_{f}$, with $A=-0.1$ using (left) only the standard Friedman equation (1) and (right) the full Friedman equation (6) including the appropriate $\dot{G}/G$ terms.
@@
Turning to deuterium (Figs. $3-4$), we see a similar effect but with an even larger magnitude. As expected, a simple step function change in $G$ in the unmodified Friedman equation results in change in (D/H) with the same sign as the change in $G$. However, making use of Eq. (6) gives an additional contribution to $\dot{a}/a$ with the opposite sign, as it does for ${}^{4}$e. In the case of deuterium, this effect is so large that it reverses the sign of the change in (D/H) over almost all of the parameter space: an initial value of $G$ larger than $G_{N}$ gives a reduction in (D/H), while an initial $G$ smaller than $G_{N}$ corresponds to an increase in (D/H).
@@
Now consider the effect of an oscillatory change in $G$ (Eq. 10). We can distinguish three regimes. For $\omega^{-1}>10^{3}$ s, the value of $G$ is roughly constant during BBN with $\dot{G}/G<H$ throughout BBN, so Eqs. (13) and (14) give the primordial ${}^{4}$e and deuterium abundances in this case, with well-known limits already derived in the literature. For $\omega^{-1}\sim 1-10^{3}$ s, we find that the element abundances are a complicated function of $A$, $\omega$, and $\phi$ in Eq. (10), with results that do not provide much insight. The case $\omega^{-1}<1$ s is more amenable to the derivation of limits on the model parameters. In this case, the first term in Eq. (6) averages to zero, and the effective density is simply increased by an amount proportional to $(\dot{G}/G)^{2}$. This density increase is constant (or at least can be taken to be constant during BBN; it must, of course decay away at some time after BBN). Expressing this effective change in density in terms of a constant $\rho_{0}$, we have
@@
where the second equality is valid for $A\ll 1$ and $\omega t\gg 1$. It is clear that the effects of an oscillating $G$ with constant oscillation frequency $\omega$ cannot be modeled using the conventional parametrization in terms of an additional effective number of relativistic degrees of freedom. Instead, the ratio of $\rho_{0}$ to the background radiation density will be a sharply increasing function of time. Hence, values of $(\dot{G}/G)^{2}$ large enough to be ruled out by the deuterium abundance will have a negligible impact on $Y_{p}$, so we need consider only the former in deriving limits on this model. (Conversely, an additional stiff component with $\rho\propto a^{-6}$ has a much larger effect effect on the ${}^{4}$e abundance than on the deuterium abundance, so that only limits on the former need be considered in that case stiff ).
@@ IV Discussion
While our investigation of the effects of a rapidly-change $G$ on BBN is far from exhaustive, it does provide some interesting new results. For models motivated by scalar-tensor theories, the effects of including the additional $\dot{G}/G$ terms in the Friedman equation are profound. In the case of a step-function potential, these terms tend to cancel the effects of an initial value of $G$ that differs from the present-day $G_{N}$; in the case of deuterium, we find that they can even reverse the sign of the change in (D/H). For a rapidly oscillating $G$ with constant oscillation frequency, the effect is similar to adding a constant vacuum energy density. This dominantly effects deuterium production, allowing us to place an upper limit on the product of the oscillation frequency and amplitude.
[…]

### v2

@@ II Models with Rapidly-Varying G
Step function models for $G$ have been explored in connection with inflation Ash and, more recently, as a possible resolution of the Hubble tension Marra ; Sakr ; Benevento ; Kable ; Alestas1 ; Alestas2 . Other observational consequences of a step function change in $G$ are explored in Refs. Alestas1 ; Alestas2 ; Paraskevas . Models of this kind are easily generated in the context of scalar-tensor theories with a step function $F(\phi)$ (see, e.g., Refs. Ash ; Paraskevas ), and they naturally asymptote to $G=G_{N}$, with $\dot{G}=0$, at late times. In general, the value of $A$ is taken to be a free parameter in these models, but models that resolve the Hubble tension typically require a change in $G$ on the order of 10%.
@@ III Effects on Big Bang Nucleosynthesis
The weak reactions drop out of thermal equilibrium at $T\sim 1$ MeV, while free neutron decay continues until $T\sim 0.1$ MeV, when rapid fusion into heavier elements occurs. At this point, nearly all of the remaining neutrons become bound into 4He, with a small fraction incorporated into deuterium, tritium, and 3He. There is also production of 7Li and 7Be, with the latter decaying into the former via electron capture at the beginning of the recombination era Sunyaev . The element abundances produced in BBN depend on the baryon/photon ratio $\eta$, which can be independently determined from the CMB. We adopt a value of $\eta=6.1\times 10^{-10}$, consistent with results from Planck Ade . This value of $\eta$ yields predicted abundances of D and 4He consistent with observations. On the other hand, the 7Li abundance predicted by BBN is well known to lie a factor $\sim 3$ above the observationally-inferred value, giving rise to the so-called “lithium problem” lithium . As it is not clear whether this problem lies in the interpretation of the lithium abundance observations or in the standard model of BBN itself, we will concentrate on 4He and deuterium production. Following standard practice, we give the primordial 4He abundance as a mass fraction, $Y_{p}$, and the deuterium abundance as a number density relative to hydrogen, (D/H).
When the change in $G$ is taken to be a constant during BBN, the effect on the 4He and deuterium abundances is well fit by FOYY
$Y_{p}=Y_{p0}(G/G_{N})^{0.357},$ (13)
@@
${\rm(D/H)}={\rm(D/H)}_{0}(G/G_{N})^{0.952},$ (14)
where the 00 subscript denotes the standard values of these abundances with an unmodified gravitational constant. Although both the 4He and deuterium abundances are increasing functions of $G$, this dependence arises for different reasons in these two cases. The 4He abundance depends mostly on the abundance of free neutrons when rapid fusion commences. An increase in $G$ causes the weak rates to freeze out at a higher temperature, when the equilibrium abundance of free neutrons is larger. Since nearly all of these neutrons end up in 4He, the result is a larger 4He abundance. However, changing $G$ also has a second effect on the 4He abundance. A larger $G$ corresponds to a smaller Hubble time at a given temperature. Thus, the temperature at which rapid fusion begins corresponds to a smaller time over which free neutron decay can occur, giving more neutrons to bind into 4He. This dual dependence on $G$ can be seen in Ref. Bambi , which presents response functions for the various element abundances as a function of a change in $G$ that is effectively a delta function of time. For $Y_{p}$, the response function is always positive, with dual peaks at $T\sim$ 1 MeV and $T\sim 0.1$ MeV, corresponding to the two effects noted above. Deuterium, on the other hand, is sensitive almost entirely to the value of $G$ when fusion begins. Given an arbitrarily long time, essentially all of the deuterium would fuse into heavier nuclei; the existence of primordial deuterium is a result of the relatively short time over which fusion can occur. Increasing $G$ decreases this time, resulting in less fusion of deuterium into heavier elements and a corresponding increase in (D/H). This is evident in the response function for deuterium given in Ref. Bambi . This response function is positive and sharply peaked near $T\sim 0.1$ MeV.
Figure 1: Change in the 4He mass fraction, $\Delta Y_{p}$, for a step-function variation in $G$ (Eqs. 7--9) as a function of $t_{i}$ and $t_{f}$, with $A=-0.1$ using (left) only the standard Friedman equation (1) and (right) the full Friedman equation (6) including the appropriate $\dot{G}/G$ terms.
@@
To calculate the primordial element abundances, we have modified PRyMordial, a Python code for BBN calculations primordial , incorporating Eq. (6) for the Friedman equation. To highlight the specific effects of rapid changes in $G$, we have also calculated the primordial element abundances using the standard Friedman equation (1) with time-varying $G$. We have chosen to present our results in terms of the change in the primordial 4He mass fraction, $\Delta Y_{p}$, and the change in the number density of deuterium relative to hydrogen, $\Delta$(D/H), due to the time variation in $G$, rather than giving the altered values of $Y_{p}$ and (D/H) themselves. This approach has the advantage that $\Delta Y_{p}$ and $\Delta$(D/H) are less sensitive to the input parameters such as the baryon-to-photon ratio, the neutron lifetime, and changes to the nuclear cross sections, because altering these parameters will tend to change both the standard-model and nonstandard values of these abundances in roughly the same way. Although we provide constraints on $\dot{G}/G$ for the rapidly-oscillating models, such constraints are subject to future changes in the observed element abundances, so we consider our results for $\Delta Y_{p}$ and $\Delta$(D/H) to be more interesting and fundamental.
@@
Consider first the step-function evolution for $G$ (Eqs. 7--9). As it is impractical to scan over the entire parameter space, and we are most interested in the effect of including the $\dot{G}/G$ terms in the modified Friedman equation, we consider only two representative values for $A$ ($-0.1$ and $+0.1$) and scan over $t_{i}$ and $t_{f}$. Our results for the 4He abundances in this case with $A=-0.1$ and $A=+0.1$ are displayed in Figs. $1-2$. First consider the effect of simply altering $G$ in the standard Friedman equation without including the $\dot{G}/G$ terms. When $t_{i}$ is large, this is simply equivalent to BBN with a different (constant) value of $G$: here $G=(0.9)(G_{N})$ (Fig. 1, left) or $G=(1.1)(G_{N})$ (Fig. 2, left). As we have already noted, 4He is most sensitive to the value of $G$ at early times; for large $t_{i}$, we simply obtain a constant increase or decrease in the 4He abundance, given by the corresponding constant change in $G$ as in Eq. (13). However, when we use the full expression for the scale factor given by Eq. (6) (Figs. 1-2, right), we see that the $\dot{G}/G$ terms strongly modify the change in the primordial 4He production, producing more complex behavior. In particular, for $A<0$, we have $\dot{G}/G>0$, so the $\dot{G}/G$ terms increase the value of $\dot{a}/a$. This partially cancels the effect of the smaller initial value of $G$, producing a smaller decrease in $Y_{p}$. The opposite effect occurs when $A=+0.1$. In this case, $\dot{G}/G$ is negative, resulting in a decrease in $\dot{a}/a$ that partially cancels the effect of the initially larger value of $G$. Thus, in both cases, the effect of including the additional $\dot{G}/G$ terms in the expression for $\dot{a}/a$ in Eq. (6) is to reduce the overall effect of the change in $G$.
@@
Figure 3: Change in D/H expressed as $\Delta$(D/H)×$\times 10^{5}$ for a step-function variation in $G$ (Eqs. 7--9) as a function of $t_{i}$ and $t_{f}$, with $A=-0.1$ using (left) only the standard Friedman equation (1) and (right) the full Friedman equation (6) including the appropriate $\dot{G}/G$ terms.
@@
Turning to deuterium (Figs. $3-4$), we see a similar effect but with an even larger magnitude. As expected, a simple step function change in $G$ in the unmodified Friedman equation results in change in (D/H) with the same sign as the change in $G$. However, making use of Eq. (6) gives an additional contribution to $\dot{a}/a$ with the opposite sign, as it does for 4He. In the case of deuterium, this effect is so large that it reverses the sign of the change in (D/H) over almost all of the parameter space: an initial value of $G$ larger than $G_{N}$ gives a reduction in (D/H), while an initial $G$ smaller than $G_{N}$ corresponds to an increase in (D/H).
@@
While we have concentrated on the deuterium and 4He abundances, we have also checked to see whether there are any parameter values that significantly reduce the predicted 7Li abundance while at the same time giving acceptable abundances for deuterium and 4He, thus serving as a possible solution to the lithium problem. We find no such parameter values for the step function change in $G$. In general, a significant reduction in 7Li is always accompanied by an unacceptably large increase in the predicted deuterium abundance. A similar effect has been seen in other proposed solutions to the lithium problem using nonstandard BBN (see, e.g., Ref. lithium ).
A full set of constraints on the allowed parameter space would require a scan over the value of $A$ as well as $t_{i}$ and $t_{f}$ and would, in any case, not provide much insight into this model. However, we can use our limited set of results to make useful qualitative statements about BBN constraints on this model. In going from a constant change in $G$ to a step-function change without including the $\dot{G}/G$ terms in the Friedman equation, it is clear from Figs. $1-2$ that the step function always produces a smaller change in 4He and deuterium than does the same constant change in $G$. This is to be expected, since the only effect of the step function in this case is to reduce the time over which $G$ differs from its present-day value. Thus, any observational limits for a particular value of $A$ will always be weaker in the step-function case than in the case where the change in $G$ is constant throughout nucleosynthesis. In going from the naive step-function model to one which correctly incorporates the additional $\dot{G}/G$ terms in the Friedman equation, these limits are weakened even further when considering 4He, since the effect of adding these additional terms tends to counteract the effect of changing the value of $G$, as we have noted. However, in the case of deuterium, the situation is more complex. In this case, the sign of the effect on the deuterium abundance can change, with the most extreme values for $\Delta$(D/H) corresponding to $t_{i}=200-500$ sec and $t_{f}=400-700$ sec. Over this range, observational limits on deuterium have the potential to place tighter constraints on a step-function variation in $G$ than on models with a constant $G$. However, this range for $t_{i}$ and $t_{f}$ represents only a small slice of parameter space for these models.
Now consider the effect of an oscillatory change in $G$ (Eq. 10). We can distinguish three regimes. For $\omega^{-1}>10^{3}$ s, the value of $G$ is roughly constant during BBN with $\dot{G}/G<H$ throughout BBN, so Eqs. (13) and (14) give the primordial 4He and deuterium abundances in this case, with well-known limits already derived in the literature. For $\omega^{-1}\sim 1-10^{3}$ s, we find that the element abundances are a complicated function of $A$, $\omega$, and $\phi$ in Eq. (10), with results that do not provide much insight. The case $\omega^{-1}<1$ s is more amenable to the derivation of limits on the model parameters. In this case, the first term in Eq. (6) averages to zero, and the effective density is simply increased by an amount proportional to $(\dot{G}/G)^{2}$. This density increase is constant (or at least can be taken to be constant during BBN; it must, of course decay away at some time after BBN). Expressing this effective change in density in terms of a constant $\rho_{0}$, we have
@@
where the second equality is valid for $A\ll 1$ and $\omega t\gg 1$, for Hubble time $t$. It is clear that the effects of an oscillating $G$ with constant oscillation frequency $\omega$ cannot be modeled using the conventional parametrization in terms of an additional effective number of relativistic degrees of freedom. Instead, the ratio of $\rho_{0}$ to the background radiation density will be a sharply increasing function of time. Hence, values of $(\dot{G}/G)^{2}$ large enough to be ruled out by the deuterium abundance will have a negligible impact on $Y_{p}$, so we need consider only the former in deriving limits on this model. (Conversely, an additional stiff component with $\rho\propto a^{-6}$ has a much larger effect effect on the 4He abundance than on the deuterium abundance, so that only limits on the former need be considered in that case stiff ).
@@ IV Discussion
While our investigation of the effects of a rapidly-changing $G$ on BBN is far from exhaustive, it does provide some interesting new results. For models motivated by scalar-tensor theories, the effects of including the additional $\dot{G}/G$ terms in the Friedman equation are profound. In the case of a step-function potential, these terms tend to cancel the effects of an initial value of $G$ that differs from the present-day $G_{N}$; in the case of deuterium, we find that they can even reverse the sign of the change in (D/H). For a rapidly oscillating $G$ with constant oscillation frequency, the effect is similar to adding a constant vacuum energy density. This dominantly effects deuterium production, allowing us to place an upper limit on the product of the oscillation frequency and amplitude.
Our main results show that models with rapidly-varying $G$ provide very distinct cosmological signatures beyond previously-considered models with adiabatic variations in $G$. It would be interesting to examine these effects during other cosmological epochs. For example, it is tempting to consider whether a rapidly-oscillating $G$ with constant oscillation frequency could mimic the effect of a cosmological constant. From Eqs. (6) and (10), a present-day value of $\Omega_{\Lambda}\approx 0.7$ can be achieved for $A\omega=5.4\times 10^{-18}$ s-1. Additionally, we require that $A\ll 1$ and $\omega t_{0}\gg 1$, where $t_{0}$ is the present-day Hubble time, in order for Eq. (16) to be valid and the first term in Eq. (6) to be ignored. While one could, in principle, choose sufficiently small $A$ and sufficiently large $\omega$ to evade all other observational constraints on time-varying $G$, it is far from clear that this could originate in any realistic model.